# Characterizing Pattern Matching and Its Limits on Compositional Task Structures

### v2

@@ 6 Data Scaling Law of Pattern Matching Behaviors
Our analysis in the previous section demonstrates that stronger functional equivalence evidence leads to better generalization. A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data? Intuitively, this requires the training set to support (strongly enough) the functional equivalence of every pair of inputs that shares the same intermediate state $b$. Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{\{1,2\}}_{D}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$. Assuming that generalization is constrained by the $k$-coverage, how should the train data size scale in the token set size $\left|{\mathcal{X}}\right|$ to achieve it completely? In practical terms, this question seeks to determine the amount of data necessary and/or sufficient to cover all ID combinations with $k$-coverage, which is crucial for understanding the data cost for pattern-matching generalization. To quantify the data cost, we establish and prove the following sample-complexity upper bound for learning a 2-Hop task (the complete statement and proof are provided in App.˜F):
@@
Consider a 2-Hop task with a token set of size $n$. For a uniformly randomly sampled train dataset $D$ of size $N$, consider a learner that generalizes within the $k$-coverage of $D$. Then, for large enough $n$, the learner achieves perfect ID generalization with high probability if $N\!\gtrsim\!n^{c}$ with $c=2.5-\frac{0.5}{k}$. In contrast, the learner (with $k\!\geq\!2$) does not achive perfect ID generalization with high probability for some 2-Hop task if $n^{2}\!\lesssim\!N\!\lesssim\!n^{c}$. Here, we ignore the polylogarithmic factors in $n$.
Theorem 6.1 presents a tight sample complexity bound $\tilde{\Theta}(n^{c})$ for pattern-matching-based generalization, showing that the training dataset required to ensure full ID generalization grows polynomially in the token set size $n$, with an exponent $c\in[2,2.5)$. To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(\left|{\mathcal{X}}\right|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data (see App.˜G for the measurement details). Fig.˜5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs. $\left|{\mathcal{X}}\right|$ across different task structures. The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions. Although we derive the theoretical bound only for 2-Hop, we observe clear power-law relationships for more complex structures as well. The higher exponents for Parallel-2-Hop ($c=2.43$) and 3-Hop ($c=2.58$) tasks suggest that extra computational steps essentially add another dimension of relationships that require robust coverage, driving the steeper power-law scaling.
@@
Overall, the results support that the data scaling law is primarily determined by data properties rather than model capacity or architectures, and additional generalization mechanisms will be required to achieve milder scaling laws on such compositional tasks. [^5] The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App. G). We note that our result aligns with the practical observation that parameter scaling does not significantly improve the multi-hop reasoning capability of LLMs (yang2024large) and the data-hungry nature of compositional tasks (lake2018generalization), suggesting that these could be partly attributed to pattern-matching behaviors. We leave further analysis of the connection between scaling behavior and pattern matching as an exciting future research direction.
@@ 7 Path Ambiguity Problem as a Failure Case of Pattern Matching
Experiments show that the path ambiguity indeed hinders both generalization on the ID test set and the interpretability of intermediate state representations, as the model now establishes functional equivalence for each $x_{2}$-conditioned equivalent pair. Fig.˜6 (Middle) shows that GPT-2 can fully generalize on the ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data. [^6] For $\left|{\mathcal{X}}\right|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times\left|{\mathcal{X}}\right|^{3}\;\approx\;61\text{k}$ distinct ID triples). Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig.˜18), and the Mamba model used in Sec.˜6 shows the same trend of generalization failure (Fig.˜19). In addition, extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96; however, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig.˜6 Right). In contrast, grouping by $x_{2}$-conditioned intermediate state $(b,x_{2})$ leads to high IICG scores, showing the formation of context-dependent state representations. This context-dependence due to path ambiguity raises an interpretability concern, as standard linear probing-based techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) would not reliably identify intermediate states when a model relies on pattern matching.
Hence, a generalization mechanism other than pattern matching will be required for a robust ID generalization on complex task structures that require the access and update of intermediate states through multiple paths (e.g., planning tasks (ruis2020benchmark; kambhampati2024position; valmeekam2023planbench)), where further characterization of this problem remains as an exciting future direction.
@@ 8 CoT Improves Data Efficiency but Path Ambiguity Persists
CoT supervision (wei2022chain; kojima2022large) dramatically improves performance on multi-step reasoning tasks. We investigate how CoT interacts with our framework and whether it can address the challenges observed in Sections 6 and 7. Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example. This substantially improves data efficiency (Fig.˜7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in the 3-Hop task, aligning with previous studies on the sample efficiency of CoT (srivastava2022beyond; kim2024transformers; wen2024sparse). The scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision. We interpret this as CoT effectively ‘flattening’ multi-hop structures into sequences of single-hop tasks, reducing the compounding data requirements of deeper compositional structures.
However, we find the path ambiguity problem persists even with CoT supervision. Despite showing improvements, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig.˜7 (Middle). IICG analysis (Right) reveals that the model’s representations remain partially context-dependent. For the 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve. In contrast, the IICG score for Non-tree task is significantly elevated at every layer with the same conditioning, suggesting the absence of disentangled state representation inside the model. We hypothesize this arises since CoT supervision does not give enough evidence that different $(x_{1},x_{2})$ pairs sharing the same $b$ should yield identical second-step outputs, as functional equivalence holds only when $x_{2}=x_{2}^{\prime}$. Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec.˜7. Our analysis may explain why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data (stechly2024chain), where we leave further analysis as future work.
@@ 9.1 Practical Implications
First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization (lake2017building). For instance, our scaling-law analysis (Theorem˜6.1) provides a quantitative explanation for the observed behavior that the data demand for generalization on multi-hop natural language data dramatically increases with the number of hops (yao-etal-2025-language-models). Such a data-hungry nature of compositional tasks is also well observed in semantic parsing (dong-lapata-2016-language), where it has been shown that the diversity of data (i.e., the same component is shown in diverse contexts) is more important than the sheer size of the dataset for better generalization (keysers2019measuring). Based on these insights, we believe our results will motivate strategic data augmentation methods that seek to maximize coverage by ensuring diverse shared contexts for functionally equivalent components (andreas2019good).
Second, our framework helps interpret various real-world failure cases observed in LLMs as originating from their reliance on pattern matching. For example, our framework aligns with the reversal curse phenomenon, which is the failure to automatically infer that ‘$A$ is a child of $B$’ by observing that ‘$B$ is a parent of $A$’ (berglund2023reversal). This is because pattern matching is fundamentally incapable of generalizing to the reversing relation without explicit functional equivalence evidence for that in the training dataset. Similarly, our framework can explain the notoriously hard problem of logical negation in LLMs (truong-etal-2023-language). A purely pattern-matching learner cannot deduce that the negation rule (i.e., if a statement ‘$p$’ is true, then ‘not $p$’ is false and vice versa) should apply to any well-formed statement, including the ones that aren’t observed, through finite observations. Finally, difficulties in solving complex planning tasks with LLMs (valmeekam2023planbench; stechly2024chain; kambhampati2024position; wang2024planning) might be partially attributed to path ambiguities, since such tasks likely require the correct tracking of intermediate states, which can be affected by various computational paths.
@@ 9.2 Towards a Taxonomy of Generalization Mechanisms
Natural language tasks possess algebraic and structural properties that differ from the random mappings we studied. For example, the same knowledge can be used in any hop of the multi-hop reasoning in practice, unlike our 2-Hop task that used $f_{1}\neq f_{2}$. In practice, a learner may reasonably harness such properties of a given task to generalize beyond the boundary of pattern-matching defined by coverage. Therefore, it is natural to ask: What generalization mechanisms enable generalization beyond the coverage boundary? While a complete answer requires future work, we outline a mechanism-based taxonomy as a starting point for a constructive categorization of distinct generalization mechanisms beyond pattern matching:
@@
  • Shared-operator generalization leverages the reuse of the same computation across positions (e.g., when $f_{1}=f_{2}$ in a two-hop task), which may be important in compositional generalization. For example, it is known that Transformers with inductive biases towards computation reuse can improve generalization on compositional tasks (csordas2021devil).
@@ 10 Conclusion
In this work, we formalized a framework for characterizing pattern matching. Our theoretical and experimental analyses yielded quantitative and predictive insights into modern neural networks’ pattern-matching behaviors, moving beyond post-hoc accounts of a model’s behavior on compositional tasks: (i) the alignment of instance-wise success with the strength of functional-equivalence evidence (Sec.˜5), (ii) the theoretical identification and empirical verification of a sharp sample complexity bound for complete ID generalization on the 2-Hop task through pattern matching (Sec.˜6), and (iii) the identification of the path ambiguity problem that impairs accuracy and interpretability even under high coverage and CoT supervision (Sec.˜7 and 8). We anticipate that future work will build on this foundation, towards a more complete and constructive understanding of compositional generalization and its failures.
@@ Acknowledgements
We thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their valuable feedback on our work.
Contents of the Appendix
@@ B.1 Dataset Construction
For the 2-Hop task, we construct two primitive functions $f_{1}:{\mathcal{X}}^{2}\rightarrow{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\rightarrow{\mathcal{X}}$ by randomly mapping from their respective domains to the codomain $\mathcal{X}$}. We create the domain by taking the Cartesian product of the token set with itself. For each function, we randomly designate a fraction $p_{\text{seen}}=0.7$ of its domain as the "seen" portion, resulting in sets $S_{f_{1}}$ and $S_{f_{2}}$.
@@
We carefully construct test sets to evaluate the model’s generalization capabilities across different coverage conditions. Our test sets contain:
@@ B.2 Training Details
For our experiments, we employ three GPT-2 model variants of increasing size: GPT-2-Small (68M parameters), GPT-2 (96M parameters), and GPT-2-XL (1.5B parameters). As shown in Tab.˜1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768. The standard GPT-2 configuration used in most experiments features 8 layers with 12 attention heads while maintaining the same hidden dimension of 768. Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600. The implementation follows the codebase from (wang2024grokked).

### v3

@@ 6 Data Scaling Law of Pattern Matching Behaviors
Our analysis in the previous section demonstrates that stronger functional equivalence evidence leads to better generalization. A natural follow-up question arises: How large should the training set be to enable full generalization on all ID test data? Intuitively, this requires the training set to support (strongly enough) the functional equivalence of every pair of inputs that shares the same intermediate state $b$. Formally, for a 2-Hop task we need $(x_{1},x_{2})\equiv^{(D,k)}_{\{1,2\}}(x^{\prime}_{1},x^{\prime}_{2})$ whenever $f_{1}(x_{1},x_{2})=f_{1}(x^{\prime}_{1},x^{\prime}_{2})$. Assuming that generalization is constrained by the $k$-coverage, how should the train data size scale in the token set size $\left|{\mathcal{X}}\right|$ to achieve it completely? In practical terms, this question seeks to determine the amount of data necessary and/or sufficient to cover all ID combinations with $k$-coverage, which is crucial for understanding the data cost for pattern-matching generalization. To quantify the data cost, we establish and prove the following sample-complexity upper bound for learning a 2-Hop task (the complete statement and proof are provided in App.˜F):
@@
Consider a 2-Hop task with a token set of size $n$. For a uniformly randomly sampled train dataset $D$ of size $N$, consider a learner that generalizes within the $k$-coverage of $D$. Then, for sufficiently large $n$, the learner can achieve perfect ID generalization with high probability if $N\!\gtrsim\!n^{c}$ with $c=2.5-\frac{0.5}{k}$.
In contrast, the learner (with $k\!\geq\!2$) cannot achieve perfect ID generalization with high probability for some 2-Hop task if $n^{2}\!\lesssim\!N\!\lesssim\!n^{c}$. Here, we ignore the polylogarithmic factors in $n$.
Theorem˜6.1 presents a tight sample complexity bound $\tilde{\Theta}(n^{c})$ for pattern-matching-based generalization, showing that the training dataset required to ensure full ID generalization grows polynomially in the token set size $n$, with an exponent $c\in[2,2.5)$. To empirically confirm this, we define a practical threshold $\hat{N}_{\mathrm{req}}$ to estimate $N_{\mathrm{req}}(\left|{\mathcal{X}}\right|,k)$, as a minimal amount of training data required to exceed ID accuracy of 0.99 within 100 epochs after reaching the same level on training data (see App.˜G for the measurement details). Fig.˜5 (Left) shows the measured power-law exponents for $\hat{N}_{\text{req}}$ vs. $\left|{\mathcal{X}}\right|$ across different task structures. The measured exponent for 2-Hop ($c=2.26$) aligns well with our theoretical predictions. Although we derive the theoretical bound only for 2-Hop, we observe clear power-law relationships for more complex structures as well. The higher exponents for Parallel-2-Hop ($c=2.43$) and 3-Hop ($c=2.58$) tasks suggest that extra computational steps essentially add another dimension of relationships that require robust coverage, driving the steeper power-law scaling.
@@
Overall, the results support that the data scaling law is primarily determined by data properties rather than model capacity or architectures, and additional generalization mechanisms will be required to achieve milder scaling laws on such compositional tasks. [^5] The observed scaling relationships are robust across different hyperparameters (weight decay and learning rate) and empirical decision criteria for $\hat{N}_{\mathrm{req}}$ (see App. G). We note that our result aligns with the practical observation that parameter scaling does not significantly improve the multi-hop reasoning capability of LLMs (Yang et al., 2024) and the data-hungry nature of compositional tasks (Lake and Baroni, 2018), suggesting that these could be partly attributed to pattern-matching behaviors. We leave further analysis of the connection between scaling behavior and pattern matching as an exciting future research direction.
@@ 7 Path Ambiguity Problem as a Failure Case of Pattern Matching
Experiments show that the path ambiguity indeed hinders both generalization on the ID test set and the interpretability of intermediate state representations, as the model now establishes functional equivalence for each $x_{2}$-conditioned equivalent pair. Fig.˜6 (Middle) shows that GPT-2 can fully generalize on the ID test set of 2-Hop task within a reasonable time with increasing data size, but fails with Non-tree task, even provided with a near-exhaustive amount of possible ID combinations as training data. [^6] For $\left|{\mathcal{X}}\right|=50$ and $p_{\text{seen}}=0.7$, our largest run ($N=50\text{k}$) includes virtually the entire domain ($\approx 0.7^{2}\!\times\left|{\mathcal{X}}\right|^{3}\;\approx\;61\text{k}$ distinct ID triples). Notably, scaling to 1.5B parameters does not show significant improvement in the performance (Fig.˜18), and the Mamba model used in Sec.˜6 shows the same trend of generalization failure (Fig.˜19). In addition, extremely prolonged training (36k epochs) with near-exhaustive ID combinations eventually achieves ID accuracy of 0.96; however, IICG analysis reveals no evidence of a unified intermediate state representation formation, with near-zero IICG scores when grouping by the intermediate state value $b$ (Fig.˜6 Right). In contrast, grouping by $x_{2}$-conditioned intermediate state $(b,x_{2})$ leads to high IICG scores, showing the formation of context-dependent state representations. This context-dependence due to path ambiguity raises an interpretability concern, as standard linear probing-based techniques like logit lens (nostalgebraist, 2020; Belrose et al., 2023) would not reliably identify intermediate states when a model relies on pattern matching.
Hence, a generalization mechanism other than pattern matching will be required for a robust ID generalization on complex task structures that require the access and update of intermediate states through multiple paths (e.g., planning tasks (Ruis et al., 2020; Kambhampati et al., 2024; Valmeekam et al., 2023)), where further characterization of this problem remains as an exciting future direction.
@@ 8 CoT Improves Data Efficiency but Path Ambiguity Persists
The CoT supervision (Wei et al., 2022; Kojima et al., 2022) dramatically improves performance on multi-step reasoning tasks. We investigate how it interacts with our framework and whether it can address the challenges observed in Sec.˜6 and 7. Specifically, we train models to sequentially generate intermediate states before final outputs, making 2-Hop a two-token prediction task: $(x_{1},x_{2},x_{3})\mapsto(b,t)$, for example. This substantially improves data efficiency (Fig.˜7 (Left)), with the power-law exponent dropping from 2.58 to 1.76 in the 3-Hop task, aligning with previous studies on the sample efficiency of CoT (Srivastava et al., 2023; Kim and Suzuki, 2025; Wen et al., 2025). The scaling exponents measured for 2-Hop, 3-Hop, and even 5-Hop tasks become nearly identical with CoT supervision. We interpret this as CoT effectively ‘flattening’ multi-hop structures into sequences of single-hop tasks, reducing the compounding data requirements of deeper compositional structures.
However, we find the path ambiguity problem persists even with CoT supervision. Despite showing improvements, the models fail to achieve perfect ID generalization under the same training conditions that yield perfect performance in 2-Hop task (Fig.˜7 (Middle). IICG analysis (Right) reveals that the model’s representations remain partially context-dependent. For the 2-Hop task, the representations cluster purely by intermediate states $b$, as indicated by the result that IICG measurement with $x_{2}$-conditioned states does not significantly shift the curve. In contrast, the IICG score for Non-tree task is significantly elevated at every layer with the same conditioning, suggesting the absence of disentangled state representation inside the model. We hypothesize this arises since CoT supervision does not give enough evidence that different $(x_{1},x_{2})$ pairs sharing the same $b$ should yield identical second-step outputs, as functional equivalence holds only when $x_{2}=x_{2}^{\prime}$. Hence, while CoT supervision helps with sequential computation by breaking down multi-hop structures, it may partially inherit the limitations on handling tasks with path ambiguities we describe in Sec.˜7. Our analysis may explain why LLMs struggle with complex planning tasks even when using CoT techniques and massive training data (Stechly et al., 2024), where we leave further analysis as future work.
@@ 9.1 Practical Implications
First, it accounts for the data-hungry nature of compositional tasks by demonstrating that robust coverage is required for reliable generalization (Lake et al., 2017). For instance, our scaling-law analysis (Theorem˜6.1) provides a quantitative explanation for the observed behavior that the data demand for generalization on multi-hop natural language data dramatically increases with the number of hops (Yao et al., 2025). Such a data-hungry nature of compositional tasks is also well observed in semantic parsing (Dong and Lapata, 2016), where it has been shown that the diversity of data (i.e., the same component is shown in diverse contexts) is more important than the sheer size of the dataset for better generalization (Keysers et al., 2020). Based on these insights, we believe our results will motivate strategic data augmentation methods that seek to maximize coverage by ensuring diverse shared contexts for functionally equivalent components (Andreas, 2020).
Second, our framework helps interpret various real-world failure cases observed in LLMs as stemming from their reliance on pattern matching. For example, our framework aligns with the reversal curse phenomenon, which is the failure to automatically infer that ‘$A$ is a child of $B$’ by observing that ‘$B$ is a parent of $A$’ (Berglund et al., 2024). This occurs because pattern matching inherently cannot generalize to the reversing relation without clear evidence of functional equivalence in the training dataset. Similarly, our framework can explain the notoriously hard problem of logical negation in LLMs (Truong et al., 2023). A purely pattern-matching learner cannot deduce that the negation rule (i.e., if a statement ‘$p$’ is true, then ‘not $p$’ is false and vice versa) should apply to any well-formed statement, including the ones that aren’t observed, through finite observations. Finally, difficulties in solving complex planning tasks with LLMs (Valmeekam et al., 2023; Stechly et al., 2024; Kambhampati et al., 2024; Wang et al., 2024b) may partly result from path ambiguities, since such tasks likely require the correct tracking of intermediate states, which can be affected by multiple computational paths.
@@ 9.2 Towards a Taxonomy of Generalization Mechanisms
Natural language tasks exhibit algebraic and structural properties that distinguish them from the randomly generated mappings we studied. For example, the same knowledge can be used in any hop of the multi-hop reasoning in practice, unlike our 2-Hop task that used $f_{1}\neq f_{2}$. In practice, a learner may reasonably harness such properties of a given task to generalize beyond the boundary of pattern-matching defined by coverage. Therefore, it is natural to ask: What generalization mechanisms enable generalization beyond the coverage boundary? While a complete answer requires future work, we outline a mechanism-based taxonomy as a starting point for a constructive categorization of distinct generalization mechanisms beyond pattern matching:
@@
  • Shared-operator generalization leverages the reuse of the same computation across positions (e.g., when $f_{1}=f_{2}$ in a two-hop task), which may be important in compositional generalization. For example, it is known that Transformers with inductive biases towards computation reuse can improve generalization on compositional tasks (Csordás et al., 2021).
@@ 10 Conclusion
In this work, we formalized a framework for characterizing pattern matching. Our theoretical and experimental analyses provided quantitative and predictive insights into modern neural networks’ pattern-matching behaviors, moving beyond post-hoc accounts of a model’s behavior on compositional tasks: (i) the alignment of instance-wise success with the strength of functional-equivalence evidence (Sec.˜5), (ii) the theoretical identification and empirical verification of a sharp sample complexity bound for complete ID generalization on the 2-Hop task through pattern matching (Sec.˜6), and (iii) the identification of the path ambiguity problem that impairs accuracy and interpretability even under high coverage and CoT supervision (Sec.˜7 and 8). We anticipate that future work will build on this foundation, towards a more complete and constructive understanding of compositional generalization and its failures.
@@ Acknowledgements
We thank Jaewon Oh, Jihoon Ha, Hyowon Cho, Yunjae Won, Seongyun Lee, and Boshi Wang for their invaluable feedback.
This work was supported by Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)). Hoyeon Chang, Jinho Park, Miyoung Ko, Hyeonbin Hwang, & Minjoon Seo acknowledge support by the InnoCORE program of the Ministry of Science and ICT(N10250156). Hanseul Cho acknowledges support by IITP grant funded by the Korean government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project).
@@ B.1 Dataset Construction
For the 2-Hop task, we construct two primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ by randomly mapping from their respective domains to the codomain $\mathcal{X}$}. We create the domain by taking the Cartesian product of the token set with itself. For each function, we randomly designate a fraction $p_{\text{seen}}=0.7$ of its domain as the "seen" portion, resulting in sets $S_{f_{1}}$ and $S_{f_{2}}$.
@@
We carefully construct test sets to evaluate the model’s generalization capabilities across coverage conditions. Our test sets contain:
@@ B.2 Training Details
For our experiments, we employ three GPT-2 model variants of increasing size: GPT-2-Small (68M parameters), GPT-2 (96M parameters), and GPT-2-XL (1.5B parameters). As shown in Tab.˜1, GPT-2-Small consists of 4 layers with 6 attention heads and a hidden dimension of 768. The standard GPT-2 configuration used in most experiments features 8 layers with 12 attention heads while maintaining the same hidden dimension of 768. Our largest model, GPT-2-XL, significantly scales up the architecture with 48 layers, 25 attention heads, and an increased hidden dimension of 1600. The implementation follows the codebase from (Wang et al., 2024a).